OpenStudy (anonymous):

Please help!!! Last HW question, and I cannot figure it out!!! Will medal if given good help and an explanation!! Find the equations of the lines through the point (3, 8) that are parallel to and perpendicular to the line y = 3x + 8. Use y as the dependent variable.

OpenStudy (anonymous):

@mathstudent55 please help me, I have an hour to submit, and have been finished with every other problem for 45 min...

OpenStudy (anonymous):

confirm your question please

OpenStudy (anonymous):

perpendicular lines have negative reciprocal slopes....y = 3x + 8 what is the slope of this line? what is the neg reciprocal of that?

OpenStudy (anonymous):

Using the equation y=3x+8 find a parallel and perpendicular line going through the point (3,8)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I know what the parallel and perpendicular lines would be of that equation, its getting the line of the equation to go through the point i cannot do

OpenStudy (anonymous):

with the slope and the given point...use point slope form to get new line y - y1 = m(x-x1)

OpenStudy (anonymous):

ok, so y-8=-1/3(x-3) ?

OpenStudy (anonymous):

y=-1/3x+9

OpenStudy (anonymous):

yes...that is the perpendicular line

OpenStudy (anonymous):

now for the parallel line..same slope plug slope and point into formula point slope

OpenStudy (anonymous):

and so for the parallel one, just do the same thing with m=3?

OpenStudy (anonymous):

yes..and the given point

OpenStudy (anonymous):

ok, thanks!! I think it would be y=3x-1

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

uw